OpenStudy (anonymous):

The area of a triangle is given by the equation b^{2}+10b=144 where b is the base of the triangle. What is the length of the base?

OpenStudy (anonymous):

\[b^{2}+10b=144\]

OpenStudy (anonymous):

since the area of the triangle is b^2 + 10b = 144, then b^2 + 10b -144 = 0 now factor so we can find possible values of b: (b + 18)(b - 8) = 0 you have two solutions here but only one will answer your question.

OpenStudy (anonymous):

Answer choices are 8, 9, 12, 16, 18

OpenStudy (anonymous):

b^2+10b-144=0 (b+18)(b-8)=0 b=-18(which is invalid as length should be in positive value) so the answer is b=8

OpenStudy (anonymous):

b(b+10)=144=8*18....b=8